# On some indices of foliations and applications

### v1

@@ 5. Multiplicity of a foliation along a divisor of separatrices
Following [17, Section 2], we define the multiplicity of $\mathcal{F}$ along any nonempty divisor $\mathcal{B}=\sum_{B}a_{B}\cdot B$ of separatrices of $\mathcal{F}$ at $p$ as follows:
@@
The notion of balanced divisor of separatrices of a foliation was introduced by Genzmer in [21], see also [23]. By convention we set $\mu_{p}(\mathcal{F},\mathcal{B})=1$ for any empty divisor $\mathcal{B}$. In particular, when $\mathcal{B}=B_{1}+\cdots+B_{r}$ is an effective primitive divisor of separatrices of the singular foliation $\mathcal{F}$ at $p$, we recover [29, Equation (2.2), p. 329] (for the reduced plane curve $C:=\cup_{i=1}^{r}B_{i}$). Hence, if we write $\mathcal{B}=\mathcal{B}_{0}-\mathcal{B}_{\infty}$ where $\mathcal{B}_{0}$ and $\mathcal{B}_{\infty}$ are effective divisors we obtain
@@
Since $\mu(\mathcal{F},\mathcal{B}_{\infty})\geq 1$, it follows that $\mu_{p}(\mathcal{F},\mathcal{B}_{0})\geq\mu_{p}(\mathcal{F},\mathcal{B}).$ The Milnor number $\mu_{p}(\mathcal{F})$ of the foliation $\mathcal{F}$ at $p$ given by the $1$-form $\omega=P(x,y)dx+Q(x,y)dy$ is defined by $\mu_{p}(\mathcal{F})=i_{p}(P,Q).$
Remember that we consider $P$ and $Q$ coprime, so $\mu_{p}(\mathcal{F})$ is a non negative integer. In [9, Theorem A] it was proved that the Milnor number of a foliation is a topological invariant.
In [16, Section 3] the $\chi$-number of the foliation $\mathcal{F}$ at $p$ was introduced as $\chi_{p}(\mathcal{F})=\left(\sum_{q\in\mathcal{I}_{p}(\mathcal{F})}\nu_{q}(\mathcal{F})\xi_{q}(\mathcal{F})\right)-\xi_{p}(\mathcal{F})$, where $\mathcal{I}_{p}(\mathcal{F})$ denotes the set of infinitely near points of $\mathcal{F}$ at $p$ (see [16, p. 7]), $\nu_{q}(\mathcal{F})$ is the algebraic multiplicity of the strict transform of $\mathcal{F}$ passing by $q$ and $\xi_{p}(\mathcal{F})$ is the tangency excess of the foliation $\mathcal{F}$ (see [16, Definition 2.3]). By [16, Proposition 4.7] we get
@@
Since $\chi_{p}(\mathcal{F})$ is a nonnegative number (see [16, Proposition 3.1] ) we have
@@
Observe that, if $\mathcal{B}$ is a primitive balanced divisor of separatrices for $\mathcal{F}$ at $p$ and $\mu_{p}(\mathcal{F},\mathcal{B})=\mu_{p}(\mathcal{F})$ then, by [16, Proposition 3.1]), $\mathcal{F}$ is of second type or has algebraic multiplicity equals one. Moreover, if $\mathcal{F}$ is of second type and $\mathcal{B}$ is a primitive balanced divisor of separatrices of $\mathcal{F}$ then $\mu_{p}(\mathcal{F},\mathcal{B})=\mu_{p}(\mathcal{F})$.
The following proposition generalizes [16, Corollary B].
@@
It follows from (14) and (13) that $\mu_{p}(\mathcal{F})=\mu_{p}(\mathcal{F},\mathcal{B}_{0})-\mu_{p}(\mathcal{F},\mathcal{B}_{\infty})+\chi_{p}(\mathcal{F})+1$. The proof follows by applying Proposition 4.1 to $\mathcal{B}_{0}$, and substituing into the last equation. Moreover, $\mu_{p}(\mathcal{F})=\tau_{p}(\mathcal{F},\mathcal{B}_{0})$ if and only if
@@
If $\mathcal{B}$ is effective, then the last part of the proposition follows since $\mu_{p}(\mathcal{B}_{0})-\tau_{p}(\mathcal{B}_{0})$ and $\chi_{p}(\mathcal{F})$ are nonnegative integers. ∎
The following corollary generalizes [19, Corollaries 3.4].
@@
Let $\mathcal{F}$ be a germ of a singular holomorphic foliation at $(\mathbb{C}^{2},p)$. Let $\mathcal{B}$ be an effective primitive balanced divisor of separatrices. If $\mu_{p}(\mathcal{F})=\tau_{p}(\mathcal{F},\mathcal{B})$ then after a suitable holomorphic coordinate transformation $\mathcal{B}$ is quasi-homogeneous and either $\mathcal{F}$ has algebraic multiplicity $1$ at $p$ or $\mathcal{F}$ is of second type.
@@
Since $\mu_{p}(\mathcal{F})=\tau_{p}(\mathcal{F},\mathcal{B})$ and $\mathcal{B}$ is effective, by Proposition 5.1 we obtain $\mu_{p}(\mathcal{B})=\tau_{p}(\mathcal{B})$. Hence, $\mathcal{B}$ belongs to its Jacobian ideal. The corollary follows from [16, Proposition 3.1 (3)] and [33, Satz p.123, (a) and (d)]. ∎
@@ 6. Proof of Theorem A
$43\mu_{p}(\mathcal{F})-4\tau_{p}(\mathcal{F},\mathcal{B}_{0})<4\left(\chi_{p}(\mathcal{F})-\mu_{p}(\mathcal{F},\mathcal{B}_{\infty})\right)+$ 4
@@
The hypothesis $\mu_{p}(\mathcal{B}_{0})\leq\mu_{p}(\mathcal{F})$ of Proposition 6.1 is essential as the following example illustrates: we consider the Suzuki’s foliation (see for instance [12, p. 75]) given by
@@
whose separatrices are $C:x=0$ and $C_{c}:y^{2}-cx^{2}-x^{3}=0$, where $c\in\mathbb{C}$. Consider the balanced divisor of separatrices $\mathcal{B}:=C+C_{0}+C_{1}-C_{2}$. After some computations we have: $517=\mu_{0}(\mathcal{B}_{0})\not\leq\mu_{0}(\mathcal{F})=$ 5, $\mu_{0}(\mathcal{F},\mathcal{B}_{\infty})=3$, $\tau_{0}(\mathcal{F},\mathcal{B}_{0})=5$, and $\chi_{0}(\mathcal{F})=0$ since $\mathcal{F}$ is a curve generalized foliation. So the left term of (16) is $\frac{5}{3}$ and Proposition 6.1 does not hold.
@@ 6.1. Proof of Theorem A
Since $\mathcal{B}$ is an effective, primitive, balanced divisor of separatrices for $\mathcal{F}$ at $p$, we have $\mathcal{B}_{\infty}=\emptyset$ and $GSV_{p}(\mathcal{F},\mathcal{B})=\Delta(\mathcal{F},\mathcal{B})$, where $\Delta(\mathcal{F},\mathcal{B})$ denotes the polar excess number of $\mathcal{F}$ with respect to $\mathcal{B}$, and this number is nonnegative because it is a sum of nonnegative numbers (see [18, Equality(3.12)]). Hence, $GSV_{p}(\mathcal{F},\mathcal{B})=\mu_{p}(\mathcal{F},\mathcal{B})-\mu_{p}(\mathcal{B})\geq 0$ by (7), which implies that $\mu_{p}(\mathcal{F},\mathcal{B})\geq\mu_{p}(\mathcal{B})$. From (15), we get $\mu_{p}(\mathcal{F})\geq\mu_{p}(\mathcal{F},\mathcal{B})\geq\mu_{p}(\mathcal{B})$. Thus, Theorem A is a consequence of Proposition 6.1. $\square$
@@ 7. Applications to global foliations
In this section, we now work with global foliations on the complex projective plane $\mathbb{P}^{2}_{\mathbb{C}}$. Let $\mathcal{F}$ be a holomorphic foliation on $\mathbb{P}^{2}_{\mathbb{C}}$. The number of points of tangency counted with multiplicities, between $\mathcal{F}$ and a non-invariant $L\subset\mathbb{P}^{2}_{\mathbb{C}}$ is the degree of the foliation, denoted by $d$.
A holomorphic foliation $\mathcal{F}$ on $\mathbb{P}^{2}_{\mathbb{C}}$ of degree $d$, is a homogeneous $1-$ form $\omega=A(x,y,z)dx+B(x,y,z)dy+C(x,y,z)dz$ with $A,B$ and $C$ homogeneous polynomials of degree $d+1$ such that $xA+yB+zC=0$.
@@ 7.1. New proof of a global result of Cerveau-Lins Neto
In [11, p. 885] Cerveau and Lins Neto stablished the following propostion. The main ingredients in their proof are the reduction of singularities of foliations and the Poincaré-Hopf formula. Here, we provide an alternative proof using our previous results and some known local formulas.
@@
where $\delta_{p}=\dim_{\mathbb{C}}\widetilde{\mathcal{O}_{p}}/\mathcal{O}_{p}$ is the colength of the local ring $\mathcal{O}_{p}$ of $(C,p)$ in its normalization. Now, it follows from the Milnor formula [31, Theorem 10.5] that
@@ 7.2. New proof of a global result of Soares
We note that $\tau_{p}(\mathcal{F},C)\geq 1$ for all singularity $p$ of $\mathcal{F}$. Indeed, there would exist a point $q\in C$ which is a singularity of $\mathcal{F}$ with $\tau_{q}(\mathcal{F},C)=0$, then we can take local coordinates centered at $q$ such that $q=(0,0)$, $\mathcal{F}$ is given by a 1-form $\omega_{q}=P(x,y)dx+Q(x,y)dy$ and $C=\{f(x,y)=0\}$. Since $\tau_{q}(\mathcal{F},C)=0$, it follows that $1\in(P,Q,f)$ so there exist $g,h,k\in\mathbb{C}\{x,y\}$ such that
@@
However, since $q\in C$ is a singularity of $\mathcal{F}$, we have $P(0,0)=Q(0,0)=f(0,0)=0$. Substituing these values into the equation (20) leads to a contradiction.
M. Soares proved the following theorem using polar varieties and characteristic classes. Here, we present a simpler proof based on the indices studied above. Recently, in [14, Theorem 4], the author presents a more general version of Soares’s theorem in higher dimensions.
@@
(Soares [35, Theorem II]) Let $C$ be an irreducible curve of degree $d_{0}>1$ that is invariant by a foliation $\mathcal{F}$ on $\mathbb{P}^{2}_{\mathbb{C}}$ of degree $d\geq 2$ with isolated singularities. Then
@@ 7.3. Bounds of the global Tjurina number of a foliation
Teissier’s Proposition [37, Chapter II, Proposition 1.2] implies that
@@
Using (23) and (24) we obtain $\mu_{p}(C)\geq(\nu_{p}(C)-1)^{2}.$ Here $i_{p}(\mathcal{P}^{df},C)$ denotes the polar intersection number of $df$ and $C$ at $p$ (see for instance [16, Section 4]).
@@
Moreover, by [34, Corollary 3.4], if $C$ is not concurrent lines, we have
@@
Furthermore, it follows from [15, Theorem 3.2] that
@@
Finally, the theorem follows by substituing (26), (27) and Remark 7.4 into (25). ∎
@@ 7.4. Alcántara–Mozo-Fernández’s conjecture
Now, we address the following conjecture posed by Alcántara–Mozo-Fernández [1, p. 16]: Does there exist a logarithmic, non-dicritical, and quasihomogeneous foliation $\mathcal{F}$ on $\mathbb{P}^{2}_{\mathbb{C}}$, with a unique singular point $p$, having an invariant reduced algebraic curve $C$ passing through $p$?
The authors indicate that they have some evidence suggesting that this type of foliation does not exist; however, they state that they do not yet have a solution to the problem. We recall that a non-dicritical foliation $\mathcal{F}$ is quasi-homogeneous at $p$, if the set of all separatrices of $\mathcal{F}$ passing by $p$ is a germ of curve given in some coordinates by a quasi-homogeneous polynomial, see [22, p. 1657].
@@
Suppose, for contradiction, that there exists a logarithmic, non-dicritical foliation $\mathcal{F}$ on $\mathbb{P}^{2}_{\mathbb{C}}$ of degree $d$, with a unique singular point $p$, having an invariant irreducible algebraic curve $C$ of degree $d_{0}\geq 2$ passing through $p$. Since $\mathcal{F}$ is logarithmic, it follows that $\mathcal{F}$ is a generalized curve foliation at $p$ and $d_{0}=d+2$ (see [18, Section 6]). Moreover, since $\mathcal{F}$ is non-dicritical at $p$, we have
@@
by [9] and [28]. Now, since $C$ is irreducible, we can apply Płoski’s inequality [32, Lemma 2.2] to $C$ at $p$, obtaining

### v2

@@ 5. Multiplicity of a foliation along a divisor of separatrices
Following [20, Section 2], we define the multiplicity of $\mathcal{F}$ along any non-empty divisor $\mathcal{B}=\sum_{B}a_{B}\cdot B$ of separatrices of $\mathcal{F}$ at $p$ as follows:
@@
The notion of balanced divisor of separatrices of a foliation was introduced by Genzmer in [24], see also [26]. By convention we set $\mu_{p}(\mathcal{F},\mathcal{B})=1$ for any empty divisor $\mathcal{B}$. In particular, when $\mathcal{B}=B_{1}+\cdots+B_{r}$ is an effective primitive divisor of separatrices of the singular foliation $\mathcal{F}$ at $p$, we recover [31, Equation (2.2), p. 329] (for the reduced plane curve $C:=\cup_{i=1}^{r}B_{i}$). Hence, if we write $\mathcal{B}=\mathcal{B}_{0}$-$\mathcal{B}_{\infty}$ where $\mathcal{B}_{0}$ and $\mathcal{B}_{\infty}$ are effective divisors we obtain
@@
Since $\mu(\mathcal{F},\mathcal{B}_{\infty})\geq 1$, it follows that $\mu_{p}(\mathcal{F},\mathcal{B}_{0})\geq\mu_{p}(\mathcal{F},\mathcal{B}).$
In [19, Section 3] the $\chi$-number of the foliation $\mathcal{F}$ at $p$ was introduced as $\chi_{p}(\mathcal{F})=\left(\sum_{q\in\mathcal{I}_{p}(\mathcal{F})}\nu_{q}(\mathcal{F})\xi_{q}(\mathcal{F})\right)-\xi_{p}(\mathcal{F})$, where $\mathcal{I}_{p}(\mathcal{F})$ denotes the set of infinitely near points of $\mathcal{F}$ at $p$ (see [19, p. 7]), $\nu_{q}(\mathcal{F})$ is the algebraic multiplicity of the strict transform of $\mathcal{F}$ passing by $q$ and $\xi_{p}(\mathcal{F})$ is the tangency excess of the foliation $\mathcal{F}$ (see [19, Definition 2.3]). By [19, Proposition 4.7] we get
@@
Since $\chi_{p}(\mathcal{F})$ is a non-negative number (see [19, Proposition 3.1] ) we have
@@
Observe that, if $\mathcal{B}$ is a primitive balanced divisor of separatrices for $\mathcal{F}$ at $p$ and $\mu_{p}(\mathcal{F},\mathcal{B})=\mu_{p}(\mathcal{F})$ then, by [19, Proposition 3.1]), $\mathcal{F}$ is of second type or has algebraic multiplicity equal to one. Moreover, if $\mathcal{F}$ is of second type and $\mathcal{B}$ is a primitive balanced divisor of separatrices of $\mathcal{F}$ then $\mu_{p}(\mathcal{F},\mathcal{B})=\mu_{p}(\mathcal{F})$.
The following proposition generalizes [19, Corollary B].
@@
It follows from (13) and (12) that $\mu_{p}(\mathcal{F})=\mu_{p}(\mathcal{F},\mathcal{B}_{0})-\mu_{p}(\mathcal{F},\mathcal{B}_{\infty})+\chi_{p}(\mathcal{F})+1$. The proof follows by applying Proposition 4.1 to $\mathcal{B}_{0}$, and substituting into the last equation. Moreover, $\mu_{p}(\mathcal{F})=\tau_{p}(\mathcal{F},\mathcal{B}_{0})$ if and only if
@@
If $\mathcal{B}$ is effective, then the last part of the proposition follows since $\mu_{p}(\mathcal{B}_{0})-\tau_{p}(\mathcal{B}_{0})$ and $\chi_{p}(\mathcal{F})$ are non-negative integers. ∎
The following corollary generalizes [22, Corollaries 3.4].
@@
Let $\mathcal{F}$ be a germ of a singular holomorphic foliation at $(\mathbb{C}^{2},p)$. Let $\mathcal{B}$ be an effective primitive balanced divisor of separatrices. If $\mu_{p}(\mathcal{F})=\tau_{p}(\mathcal{F},\mathcal{B})$ then, after a suitable holomorphic coordinate transformation, $\mathcal{B}$ is quasi-homogeneous and either $\mathcal{F}$ has algebraic multiplicity $1$ at $p$ or $\mathcal{F}$ is of second type.
@@
Since $\mu_{p}(\mathcal{F})=\tau_{p}(\mathcal{F},\mathcal{B})$ and $\mathcal{B}$ is effective, by Proposition 5.1 we obtain $\mu_{p}(\mathcal{B})=\tau_{p}(\mathcal{B})$. Hence, $\mathcal{B}$ belongs to its Jacobian ideal. The corollary follows from [19, Proposition 3.1 (3)] and [36, Satz p.123, (a) and (d)]. ∎
@@ 6. Proof of Theorem A
$3\mu_{p}(\mathcal{F})-4\tau_{p}(\mathcal{F},\mathcal{B}_{0})<4\left(\chi_{p}(\mathcal{F})-\mu_{p}(\mathcal{F},\mathcal{B}_{\infty})\right)+4$
@@
The hypothesis $\mu_{p}(\mathcal{B}_{0})\leq\mu_{p}(\mathcal{F})$ of Proposition 6.1 is essential as the following example illustrates: we consider the Suzuki’s foliation (see for instance [13, p. 75]) given by
@@
whose separatrices are $C:x=0$ and $C_{c}:y^{2}-cx^{2}-x^{3}=0$, where $c\in\mathbb{C}$. Consider the balanced divisor of separatrices $\mathcal{B}:=C+C_{0}+C_{1}-C_{2}$. After some computations we have: $17=\mu_{0}(\mathcal{B}_{0})\not\leq\mu_{0}(\mathcal{F})=5$, $\mu_{0}(\mathcal{F},\mathcal{B}_{\infty})=3$, $\tau_{0}(\mathcal{F},\mathcal{B}_{0})=5$, and $\chi_{0}(\mathcal{F})=0$ since $\mathcal{F}$ is a curve generalized foliation. So the left term of (15) is $\frac{5}{3}$ and Proposition 6.1 does not hold.
@@ 6.1. Proof of Theorem A
Since $\mathcal{B}$ is an effective, primitive, balanced divisor of separatrices for $\mathcal{F}$ at $p$, we have $\mathcal{B}_{\infty}=\emptyset$ and $GSV_{p}(\mathcal{F},\mathcal{B})=\Delta_{p}(\mathcal{F},\mathcal{B})$, where $\Delta_{p}(\mathcal{F},\mathcal{B})$ denotes the polar excess number of $\mathcal{F}$ with respect to $\mathcal{B}$, and this number is non-negative because it is a sum of non-negative numbers (see [21, Equality(3.12)]). Hence, $GSV_{p}(\mathcal{F},\mathcal{B})=\mu_{p}(\mathcal{F},\mathcal{B})-\mu_{p}(\mathcal{B})\geq 0$ by (7), which implies that $\mu_{p}(\mathcal{F},\mathcal{B})\geq\mu_{p}(\mathcal{B})$. From (14), we get $\mu_{p}(\mathcal{F})\geq\mu_{p}(\mathcal{F},\mathcal{B})\geq\mu_{p}(\mathcal{B})$. Thus, Theorem A is a consequence of Proposition 6.1. $\square$
@@ 7. Applications to global foliations
In this section, we work with global foliations on the complex projective plane $\mathbb{P}^{2}_{\mathbb{C}}$. Let $\mathcal{F}$ be a holomorphic foliation on $\mathbb{P}^{2}_{\mathbb{C}}$. The number of points of tangency counted with multiplicities, between $\mathcal{F}$ and a non-invariant $L\subset\mathbb{P}^{2}_{\mathbb{C}}$ is the degree of the foliation, denoted by $d$.
A holomorphic foliation $\mathcal{F}$ on $\mathbb{P}^{2}_{\mathbb{C}}$ of degree $d$, is a homogeneous 11- form $\omega=A(x,y,z)dx+B(x,y,z)dy+C(x,y,z)dz$ with $A,B$ and $C$ homogeneous polynomials of degree $d+1$ such that $xA+yB+zC=0$.
@@ 7.1. New proof of a global result of Cerveau-Lins Neto
In [12, p. 885] Cerveau and Lins Neto established the following proposition. The main ingredients in their proof are the reduction of singularities of foliations and the Poincaré-Hopf formula. Here, we provide an alternative proof using our previous results and some known local formulas.
@@
where $\delta_{p}=\dim_{\mathbb{C}}\widetilde{\mathcal{O}_{p}}/\mathcal{O}_{p}$ is the co-length of the local ring $\mathcal{O}_{p}$ of $(C,p)$ in its normalization. It follows from the Milnor formula [34, Theorem 10.5] that
@@ 7.2. New proof of a global result of Soares
We note that $\tau_{p}(\mathcal{F},C)\geq 1$ for all singularity $p$ of $\mathcal{F}$. Indeed, there would exist a point $q\in C$ which is a singularity of $\mathcal{F}$ with $\tau_{q}(\mathcal{F},C)=0$, then, we can take local coordinates centered at $q$ such that $q=(0,0)$, $\mathcal{F}$ is given by a 1-form $\omega_{q}=P(x,y)dx+Q(x,y)dy$ and $C=\{f(x,y)=0\}$. Since $\tau_{q}(\mathcal{F},C)=0$, it follows that $1\in(P,Q,f)$ so there exist $g,h,k\in\mathbb{C}\{x,y\}$ such that
@@
However, since $q\in C$ is a singularity of $\mathcal{F}$, we have $P(0,0)=Q(0,0)=f(0,0)=0$. Substituting these values into equation (19) leads to a contradiction.
M. Soares proved the following theorem using polar varieties and characteristic classes. In [17], the author obtained a characterization of Soares’s formula (in higher dimensions), in terms of the Schwartz and GSV indices of the foliation. Also, in [16], the author obtained another generalization of Soares’s theorem for foliations on hypersurfaces with isolated singularities. In [15], the author presents a more general version of Soares’s theorem in higher dimensions. Here, we present a simpler proof based on the indices studied above.
@@
(Soares [39, Theorem II]) Let $C$ be an irreducible curve of degree $d_{0}>1$ that is invariant by a foliation $\mathcal{F}$ on $\mathbb{P}^{2}_{\mathbb{C}}$ of degree $d\geq 2$ with isolated singularities. Then
@@ 7.3. Bounds of the global Tjurina number of a foliation
Teissier’s Proposition [41, Chapter II, Proposition 1.2] implies that
@@
Using (22) and (23) we obtain $\mu_{p}(C)\geq(\nu_{p}(C)-1)^{2}.$
@@
Moreover, by [38, Corollary 3.4], if $C$ is not concurrent lines, we have
@@
Furthermore, it follows from [18, Theorem 3.2] that
@@
Finally, the theorem follows by substituting (25), (26) and Remark 7.4 into (24). ∎
@@ 7.4. Alcántara–Mozo-Fernández’s conjecture
Now, we address the following conjecture posed by Alcántara–Mozo-Fernández [1, p. 16]: Does there exist a logarithmic, non-dicritical, and quasi-homogeneous foliation $\mathcal{F}$ on $\mathbb{P}^{2}_{\mathbb{C}}$, with a unique singular point $p$, having an invariant reduced algebraic curve $C$ passing through $p$?
These authors indicate that they have some evidence suggesting that this type of foliation does not exist; however, they state that they do not yet have a solution to the problem.
Recall that a foliation $\mathcal{F}$ on $\mathbb{P}^{2}_{\mathbb{C}}$ is logarithmic if there exist homogeneous, irreducible polynomials $f_{1},\ldots,f_{m}\in\mathbb{C}[x,y,z]$, of degrees $d_{1},\ldots,d_{m}$, respectively, such that a 1-form describing the foliation is:
$\Omega=f_{1}\cdot\ldots\cdot f_{m}\sum_{i=1}^{m}\lambda_{i}\frac{df_{i}}{f_{i}},$
for some non-zero complex numbers $\lambda_{1},\ldots,\lambda_{m}$ satisfying $\sum_{i=1}^{m}\lambda_{i}d_{i}=0$. Note that the foliation has degree $\deg(\mathcal{F})=\sum_{i=1}^{m}d_{i}-2$. In particular, taking $C=\{f_{1}\cdots f_{m}=0\}$, we obtain
(27) $\deg(C)=\deg(\mathcal{F})+2.$
A logarithmic foliation $\mathcal{F}$ on $\mathbb{P}^{2}_{\mathbb{C}}$ is always a generalized curve foliation, by [21, Section 6].
We also recall that a non-dicritical foliation $\mathcal{F}$ is quasi-homogeneous at $p$, if the set of all separatrices of $\mathcal{F}$ passing by $p$ is a germ of curve given in some coordinates by a quasi-homogeneous polynomial, see [25, p. 1657].
@@
Suppose, for contradiction, that there exists a logarithmic, non-dicritical foliation $\mathcal{F}$ on $\mathbb{P}^{2}_{\mathbb{C}}$ of degree $d$, with a unique singular point $p$, having an invariant irreducible algebraic curve $C$ of degree $d_{0}\geq 2$ passing through $p$. Since $\mathcal{F}$ is logarithmic, it follows that $\mathcal{F}$ is a generalized curve foliation at $p$ and $d_{0}=d+2$ (see [21, Section 6]). Moreover, since $\mathcal{F}$ is non-dicritical at $p$, we have
@@
by [9, Theorem 4] and [7, p. 19]. Now, since $C$ is irreducible, we can apply Płoski’s inequality [35, Lemma 2.2] to $C$ at $p$, obtaining